a butcher mixes 12 lb of ground pork at $1.25 per pound with 24 lb of ground beef at $2 per pound to sell as meat loaf mix. What should be the cost per pound of the mixture
Answer by nerdybill(7384)

a butcher mixes 12 lb of ground pork at $1.25 per pound with 24 lb of ground beef at $2 per pound to sell as meat loaf mix. What should be the cost per pound of the mixture
.
Let x = cost ($) per pound of mixture
then
12(1.25) + 24(2) = x(12+24)
12(1.25) + 24(2) = x(36)
dividing both sides by 2:
6(1.25) + 24 = x(18)
7.50 + 24 = x(18)
7.50 + 24 = x(18)
31.50 = x(18)
$1.75 = x